Sweeten Company had no jobs in progress at the beginning of March and no beginning inventories. The company has two manufacturing departments--Molding and Fabrication. It started, completed, and sold only two jobs during March—Job P and Job Q. The following additional information is available for the company as a whole and for Jobs P and Q (all data and questions relate to the month of March): Molding Fabrication Total Estimated total machine-hours used 2,500 1,500 4,000 Estimated total fixed manufacturing overhead $ 14,250 $ 17,550 $ 31,800 Estimated variable manufacturing overhead per machine-hour $ 3.10 $ 3.90 Job P Job Q Direct materials $ 30,000 $ 16,500 Direct labor cost $ 34,600 $ 14,300 Actual machine-hours used: Molding 3,400 2,500 Fabrication 2,300 2,600 Total 5,700 5,100 Sweeten Company had no underapplied or overapplied manufacturing overhead costs during the month. Required: For questions 1-8, assume that Sweeten Company uses a plantwide predetermined overhead rate with machine-hours as the allocation base. For questions 9-15, assume that the company uses departmental predetermined overhead rates with machine-hours as the allocation base in both departments. Foundational 2-2 2. How much manufacturing overhead was applied to Job P and how much was applied to Job Q?
Answer #1

Solution:

Total estimated overhead = $31800 + 3.10*2500 + 3.90*1500 = $45,400

Predetrmined overhead rate = $45400/ 4000 = $11.35 per machine hour

Overhead applied to Job P = $11.35*5700 = $64,695

Overhead applidd to Job Q = $11.35*5100 = $57,885